# Mamba-Driven Topology Fusion for Monocular 3-D Human Pose Estimation

### v1

@@ III-D Bone-Joint Fusion Embedding
Figure 4: The structural diagrams of the spatiotemporal refinement module and the bone-joint fusion embedding. The bone-joint fusion embedding serves as the head of the spatiotemporal refinement module, processing the input 2-D pose sequences and 3-D bone sequences.
Before feeding the 2-D pose and bone vectors into the refinement network, it is necessary to fuse information from the two representations. As they reside in different coordinate systems, we first map both representations into high-dimensional feature spaces, extract salient features, and then perform feature fusion. The structure of the bone-joint fusion embedding is illustrated in Figure 4. For the 2-D pose sequence branch, the 2-D sequence is first mapped to high-dimensional embedding features, followed by adding joints positional embedding $E_{s}\in\mathbb{R}^{j\times D}$, and then processed by a GEM block:
@@
For the bone vector branch, we use the same processing approach as the 2-D pose branch, but without adding positional embedding. Instead, we use VIM to process spatial relationships, as the graph convolutional structure in GEM is designed for joints. This can be expressed as follows:
@@
Since 2-D poses and bone vectors are two different attributes, the features processed through spatiotemporal extraction are more conducive to the fusion of these two types of information. Ultimately, we adopt an adaptive fusion method [19] to aggregate both sequence information, which is defined as:
@@ III-E Spatiotemporal Refinement Module
To predict the final 3-D pose sequence $s_{3D}$, $X_{0}$ is sent into the spatiotemporal refinement module. As shown in Figure 4, the spatiotemporal refinement module consists of multiple layers of basic blocks, each layer comprising a series of GEM and VIM blocks. The computation process for the $i$-th layer can be expressed as:
@@
where $i\in\{0,1,\dots,L\}$, $X_{i+1}\in\mathbb{R}^{f\times j\times D}$. GEM is used to handle the relationships between local and global keypoints in the spatial pose sequence, while VIM is used to capture the relationships of the same keypoint across consecutive time frames in the temporal pose sequence. After multiple layers of spatiotemporal refinement, a regression head is used to compute the precise 3-D pose sequence $s_{3D}\in\mathbb{R}^{f\times j\times 3}$.
@@ IV-A Dataset and evaluation metric
We evaluate our method on two 3-D HPE bench-mark datasets: Human3.6M [34] and MPI-INF-3DHP [35].
@@
The Human3.6M dataset consists of 3.6 million video frames captured in indoor settings, featuring 11 professional actors performing 15 different actions across 4 synchronized camera views. For training and testing, we follow the approach in [25, 40], using data from 5 subjects (S1, S5, S6, S7, S8) for training and 2 subjects (S9, S11) for testing. We evaluate model performance using two protocols, as in prior work [11, 40]. Protocol 1 calculates the mean per-joint position error (MPJPE), which measures the average Euclidean distance in millimeters between predicted and ground-truth 3-D keypoints. Protocol 2 uses procrustes MPJPE (P-MPJPE), which computes the reconstruction error after aligning the predicted 3-D pose to the ground-truth using procrustes analysis. Moreover, following Pavllo et al. [40], we evaluate the model’s MPJVE, which is the mean Euclidean distance of the first derivative.
@@ IV-A2 MPI-INF-3DHP
The MPI-INF-3DHP dataset, another popular benchmark for 3-D human pose estimation, includes 1.3 million frames from both controlled indoor environments and complex outdoor scenarios. Following the setup in [11, 41], we use MPJPE, the percentage of correct keypoints (PCK) with a 150mm threshold, and the area under the curve (AUC) as evaluation metrics.
@@ IV-B Implementation Details
The experiments are conducted using PyTorch and run on an RTX 4090 GPU. 2-D pose sequences are obtained by the Stacked Hourglass (SH) networks [12]. The training process is divided into two stages. In the first stage, the bone aware module is trained using a cross-entropy loss function to optimize the model. After the training of the first stage is completed, the parameters of the bone perception module are frozen. In the second stage, only the spatiotemporal refinement module is trained, using weighted MPJPE, N-MPJPE, and MPJVE as the loss functions, with weights of 1, 0.5, and 20, respectively. The learning rate, learning rate decay, and dropout for the first stage are set to $2\times 10^{-3}$, $0.99$, and $0.1$, respectively. For the second stage, the learning rate, learning rate decay, and dropout are set to $5\times 10^{-4}$, $0.99$, and $0.1$, respectively. Both modules are trained using the AdamW optimizer [55] with weight decay of $0.01$. The bone aware module is trained for 60 epochs with a batch size of 128, and the spatiotemporal refinement module is trained for 120 epochs with a batch size of 16.
@@ IV-C Quantitative Evaluation
Ours-large 243 4.4M 58.7M 40.0mm
[…]
@@
TABLE III: MPJVE Comparison for Each Action on Human3.6M Using Detected 2-D Pose Sequences.
@@
As shown in Table II, our method outperforms previous SOTA methods on both MPJPE and P-MPJPE (40.0 mm on Protocol #1 and 33.5 mm on Protocol #2). Our method performs better on most actions, especially on challenging actions with significant occlusions, such as sitting and sitting down. This indicates that our proposed topology enhanced mamba can effectively capture both local and global topological structures of the keypoints. To explore the upper bound of model performance, results using real 2-D pose inputs are reported in the middle of Table II. Our method achieves the best results on all actions compared to previous methods (15.4 mm on Protocol #1). This shows that using better 2-D pose estimation results as input can amplify the expressive power of our model.
@@ IV-C2 Results on MPI-INF-3DHP
To evaluate the generalization capability of the model, we test on the MPI-INF-3DHP dataset using the MPJPE, PCK, and AUC metrics. Real 2-D pose estimates are used as input. As shown in Table IV, the Ours-large method achieves the best results in PCK, AUC, and MPJPE metrics, surpassing SOTA models with lower computational cost. The Ours-tiny method also achieves competitive results in MPJPE with an extremely small computational budget. These findings demonstrate that our approach is robust and performs well in both indoor and outdoor scenarios.
@@ IV-D Qualitative Analysis
Figure 6: Visualization Comparison of Ours-large and PoseMamba Using 2-D Poses with Erroneous Estimations.
@@
Figure 6 illustrates a comparison of joint positions between Ours-large and PoseMamba-B  [28] on actions with significant occlusions. When the input 2-D poses contain substantial errors, our method demonstrates greater robustness compared to PoseMamba-B.
@@ IV-E1 Analysis on All Modules
VIM + GEM(Spatiotemporal Refinement Module) 43.0 36.0
@@
Figure 8: The comparison of computational cost (upper) and MPJPE (lower) between Ours-Tiny and StridedFormer with respect to different input sequence lengths. In the MACs comparison, $m\times$ indicates the multiplier of the computational cost for the current sequence length relative to that of a 27-frame sequence. For example, 10.7×10.7$\times$10.7 × implies that the computational cost of the 243-frame model ($686M$) is 10.7 times that of the 27-frame model ($64M$).
We evaluated the effectiveness of various modules, including the VIM, GEM, bone aware Module, and bone-joint fusion embedding, as presented in Table V. The baseline model replaces the basic blocks in the spatiotemporal refinement module with Transformer encoder blocks, setting the depth to 8 and the hidden state dimension to 64, while keeping the remaining parameter settings consistent with the spatiotemporal refinement module. It can be observed that when the Transformer blocks are replaced with VIM blocks of the same configuration, the error decreases significantly. This is because previous Transformer-based pose estimators heavily relied on high-dimensional features, whereas we have set the feature dimension to 64. In contrast, VIM blocks are more suitable for low-dimensional features, significantly reducing computational cost while maintaining superior accuracy compared to Transformer blocks (44.0 mm vs. 59.2 mm). When the VIM blocks for processing spatial features are replaced with GEM blocks, the model (i.e., our spatiotemporal refinement module) achieves a further 1.0 mm reduction in MPJPE. This demonstrates that, building on the modeling capability of the Mamba model, our convolutional structure enhanced with graph convolution more effectively captures the spatial dependencies among joints. By directly concatenating the outputs of the bone aware module with the 2-D pose and projecting them into a high-dimensional space through a linear layer, the MPJPE only decreases by 0.5 mm. However, employing the proposed bone-joint fusion embedding to learn representations from both coordinate systems before fusion proves more effective in guiding subsequent models to capture joint positions, resulting in a 1.3mm reduction in MPJPE.
@@ IV-E2 Bone Aware Module
TABLE VII: The network architecture of the Bone Aware Module is illustrated, where $Depth$ denotes the network depth, $Dim$ the hidden feature dimension.
@@
The ablation study on the bone aware module is using detected 2-D poses as input. First, we analyze the impact of the number of predicted polar angle categories in the bone aware module. As shown in the upper section of Table VI, an ablation study is performed across various category settings. When the classification is limited to two categories, the bone aware module achieves the highest classification accuracy of 91%; however, this configuration also results in the largest error. This is because classifying into only two categories provides only two directional cues for bone depth, which is overly coarse and fails to offer meaningful guidance for the model. As the number of categories increases, the bone aware module provides finer directional guidance, but the classification accuracy decreases accordingly. The best MPJPE is achieved with six categories. Additionally, we evaluate the coordinate representation of the output bone vectors. As shown in the lower section of Table VI, spherical coordinates perform better than Cartesian coordinates because they incorporate explicit information about bone angles and lengths.
@@ IV-E3 Spatiotemporal Refinement Module
TABLE VIII: Ablation study on the fusion positions of GCN and Mamba modules. $Parallel$ denotes parallel processing streams of GCN and Mamba. $Sequential(GCN+Mamba)$ denotes the configuration where the input is first processed by GCN followed by Mamba, whereas $Sequential(Mamba+GCN)$ denotes the reverse order. $Inner$ indicates that GCN is embedded within the internal structure of the Mamba block.
@@
[…]
@@
[…]
@@ V Conclusion
In this paper, we propose a Mamba-Driven Topology Fusion framework for monocular 3-D human pose estimation. We first design a bone aware module to estimate the direction and position of bone vectors, providing topological guidance for the Mamba model in capturing dependencies within joint sequences. Based on the Vision Mamba block, we further introduce GEM, which incorporates internal forward and backward Graph Convolutional Networks to enhance Mamba’s ability to capture local joint relationships. Finally, we propose a bone-joint fusion embedding to integrate joint and bone features across different coordinate systems, along with a spatiotemporal refinement module for accurate 3-D human pose regression. Our method achieves superior results on the Human3.6M and MPI-INF-3DHP datasets compared to previous methods, achieving a win-win in both speed and accuracy.

### v2

@@ III-D Bone-Joint Fusion Embedding
Figure 4: The structural diagrams of the spatiotemporal refinement module and the bone-joint fusion embedding. The bone-joint fusion embedding serves as the head of the spatiotemporal refinement module, processing the input 2D pose sequences and 3D bone sequences.
Before feeding the 2D pose and bone vectors into the refinement network, it is necessary to fuse information from the two representations. As they reside in different coordinate systems, we first map both representations into high-dimensional feature spaces, extract salient features, and then perform feature fusion. The structure of the bone-joint fusion embedding is illustrated in Figure 4. For the 2D pose sequence branch, the 2D sequence is first mapped to high-dimensional embedding features, followed by adding joints positional embedding $E_{s}\in\mathbb{R}^{j\times D}$, and then processed by a GEM block:
@@
For the bone vector branch, we use the same processing approach as the 2D pose branch, but without adding positional embedding. Instead, we use VIM to process spatial relationships, as the graph convolutional structure in GEM is designed for joints. This can be expressed as follows:
@@
Since 2D poses and bone vectors are two different attributes, the features processed through spatiotemporal extraction are more conducive to the fusion of these two types of information. Ultimately, we adopt an adaptive fusion method [19] to aggregate both sequence information, which is defined as:
@@ III-E Spatiotemporal Refinement Module
To predict the final 3D pose sequence $s_{3D}$, $X_{0}$ is sent into the spatiotemporal refinement module. As shown in Figure 4, the spatiotemporal refinement module consists of multiple layers of basic blocks, each layer comprising a series of GEM and VIM blocks. The computation process for the $i$-th layer can be expressed as:
@@
where $i\in\{0,1,\dots,L\}$, $X_{i+1}\in\mathbb{R}^{f\times j\times D}$. GEM is used to handle the relationships between local and global keypoints in the spatial pose sequence, while VIM is used to capture the relationships of the same keypoint across consecutive time frames in the temporal pose sequence. After multiple layers of spatiotemporal refinement, a regression head is used to compute the precise 3D pose sequence $s_{3D}\in\mathbb{R}^{f\times j\times 3}$.
@@ IV-A Dataset and evaluation metric
We evaluate our method on two 3D HPE bench-mark datasets: Human3.6M [34] and MPI-INF-3DHP [35].
@@
The Human3.6M dataset consists of 3.6 million video frames captured in indoor settings, featuring 11 professional actors performing 15 different actions across 4 synchronized camera views. For training and testing, we follow the approach in [25, 40], using data from 5 subjects (S1, S5, S6, S7, S8) for training and 2 subjects (S9, S11) for testing. We evaluate model performance using two protocols, as in prior work [11, 40]. Protocol 1 calculates the mean per-joint position error (MPJPE), which measures the average Euclidean distance in millimeters between predicted and ground-truth 3D keypoints. Protocol 2 uses procrustes MPJPE (P-MPJPE), which computes the reconstruction error after aligning the predicted 3D pose to the ground-truth using procrustes analysis. Moreover, following Pavllo et al. [40], we evaluate the model’s MPJVE, which is the mean Euclidean distance of the first derivative.
@@ IV-A2 MPI-INF-3DHP
The MPI-INF-3DHP dataset, another popular benchmark for 3D human pose estimation, includes 1.3 million frames from both controlled indoor environments and complex outdoor scenarios. Following the setup in [11, 41], we use MPJPE, the percentage of correct keypoints (PCK) with a 150mm threshold, and the area under the curve (AUC) as evaluation metrics.
@@ IV-B Implementation Details
The experiments are conducted using PyTorch and run on an RTX 4090 GPU. 2D pose sequences are obtained by the Stacked Hourglass (SH) networks [12]. The training process is divided into two stages. In the first stage, the bone aware module is trained using a cross-entropy loss function to optimize the model. After the training of the first stage is completed, the parameters of the bone perception module are frozen. In the second stage, only the spatiotemporal refinement module is trained, using weighted MPJPE, N-MPJPE, and MPJVE as the loss functions, with weights of 1, 0.5, and 20, respectively. The learning rate, learning rate decay, and dropout for the first stage are set to $2\times 10^{-3}$, $0.99$, and $0.1$, respectively. For the second stage, the learning rate, learning rate decay, and dropout are set to $5\times 10^{-4}$, $0.99$, and $0.1$, respectively. Both modules are trained using the AdamW optimizer [55] with weight decay of $0.01$. The bone aware module is trained for 60 epochs with a batch size of 128, and the spatiotemporal refinement module is trained for 120 epochs with a batch size of 16.
@@ IV-C Quantitative Evaluation
Ours-large 243 4.4M 60.7M 40.0mm
TABLE II: Quantitative comparisons for each action on Human3.6M using detected 2D pose sequence and 2D ground truth.TT represents the input time frame. (*) denotes using HRNet [13]. $\dagger$} indicates the use of 2D ground truth as inputs. The best results are highlighted in bold, and the second-best results are underlined.
@@
TABLE III: MPJVE Comparison for Each Action on Human3.6M Using Detected 2D Pose Sequences.
@@
As shown in Table II, our method outperforms previous SOTA methods on both MPJPE and P-MPJPE (40.0 mm on Protocol #1 and 33.5 mm on Protocol #2). Our method performs better on most actions, especially on challenging actions with significant occlusions, such as sitting and sitting down. This indicates that our proposed topology enhanced mamba can effectively capture both local and global topological structures of the keypoints. To explore the upper bound of model performance, results using real 2D pose inputs are reported in the middle of Table II. Our method achieves the best results on all actions compared to previous methods (15.4 mm on Protocol #1). This shows that using better 2D pose estimation results as input can amplify the expressive power of our model.
@@ IV-C2 Results on MPI-INF-3DHP
To evaluate the generalization capability of the model, we test on the MPI-INF-3DHP dataset using the MPJPE, PCK, and AUC metrics. Real 2D pose estimates are used as input. As shown in Table IV, the Ours-large method achieves the best results in PCK, AUC, and MPJPE metrics, surpassing SOTA models with lower computational cost. The Ours-tiny method also achieves competitive results in MPJPE with an extremely small computational budget. These findings demonstrate that our approach is robust and performs well in both indoor and outdoor scenarios.
@@ IV-D Qualitative Analysis
Figure 6: Visualization Comparison of Ours-large and PoseMamba Using 2D Poses with Erroneous Estimations.
@@
Figure 6 illustrates a comparison of joint positions between Ours-large and PoseMamba-B  [28] on actions with significant occlusions. When the input 2D poses contain substantial errors, our method demonstrates greater robustness compared to PoseMamba-B.
@@ IV-E1 Analysis on All Modules
VIM + GEM(Spatiotemporal Refinement Module) 42.9 36.0
@@
Figure 8: The comparison of computational cost (upper) and MPJPE (lower) between Ours-Tiny and StridedFormer with respect to different input sequence lengths. In the MACs comparison, $m\times$ indicates the multiplier of the computational cost for the current sequence length relative to that of a 27-frame sequence. For example, $10.7\times$ implies that the computational cost of the 243-frame model ($686M$) is 10.7 times that of the 27-frame model ($64M$).
We evaluated the effectiveness of various modules, including the VIM, GEM, bone aware Module, and bone-joint fusion embedding, as presented in Table V. The baseline model replaces the basic blocks in the spatiotemporal refinement module with Transformer encoder blocks, setting the depth to 8 and the hidden state dimension to 64, while keeping the remaining parameter settings consistent with the spatiotemporal refinement module. It can be observed that when the Transformer blocks are replaced with VIM blocks of the same configuration, the error decreases significantly. This is because previous Transformer-based pose estimators heavily relied on high-dimensional features, whereas we have set the feature dimension to 64. In contrast, VIM blocks are more suitable for low-dimensional features, significantly reducing computational cost while maintaining superior accuracy compared to Transformer blocks (44.0 mm vs. 59.2 mm). When the VIM blocks for processing spatial features are replaced with GEM blocks, the model (i.e., our spatiotemporal refinement module) achieves a further 1.0 mm reduction in MPJPE. This demonstrates that, building on the modeling capability of the Mamba model, our convolutional structure enhanced with graph convolution more effectively captures the spatial dependencies among joints. By directly concatenating the outputs of the bone aware module with the 2D pose and projecting them into a high-dimensional space through a linear layer, the MPJPE only decreases by 0.5 mm. However, employing the proposed bone-joint fusion embedding to learn representations from both coordinate systems before fusion proves more effective in guiding subsequent models to capture joint positions, resulting in a 1.3mm reduction in MPJPE.
@@ IV-E2 Bone Aware Module
TABLE VII: The network architecture of the Bone Aware Module is illustrated, where DepthDepth denotes the network depth, DimDim the hidden feature dimension.
@@
The ablation study on the bone aware module is using detected 2D poses as input. First, we analyze the impact of the number of predicted polar angle categories in the bone aware module. As shown in the upper section of Table VI, an ablation study is performed across various category settings. When the classification is limited to two categories, the bone aware module achieves the highest classification accuracy of 91%; however, this configuration also results in the largest error. This is because classifying into only two categories provides only two directional cues for bone depth, which is overly coarse and fails to offer meaningful guidance for the model. As the number of categories increases, the bone aware module provides finer directional guidance, but the classification accuracy decreases accordingly. The best MPJPE is achieved with six categories. Additionally, we evaluate the coordinate representation of the output bone vectors. As shown in the lower section of Table VI, spherical coordinates perform better than Cartesian coordinates because they incorporate explicit information about bone angles and lengths.
@@ IV-E3 Spatiotemporal Refinement Module
TABLE VIII: Ablation study on the fusion positions of GCN and Mamba modules. ParallelParallel denotes parallel processing streams of GCN and Mamba. $Sequential(GCN+Mamba)$ denotes the configuration where the input is first processed by GCN followed by Mamba, whereas $Sequential(Mamba+GCN)$ denotes the reverse order. InnerInner indicates that GCN is embedded within the internal structure of the Mamba block.
@@
TABLE IX: Ablation study on scanning direction and basic blocks in the spatiotemporal refinement module. To verify the effectiveness of bidirectional GEM under similar computational costs, $\dagger$} indicates an increase in model scale.
@@
We also conducted ablation experiments on the directions of VIM and GEM within the spatiotemporal refinement module. The results are reported in Table IX. When only the VIM block is used with an undirected scanning pose sequence, the error reaches a maximum of 44.2 mm. By adopting a bidirectional scanning approach, the error decreases to 44.0 mm. Replacing the VIM module with GEM to handle the spatial joint relationships reduces the MPJPE to 43.4 mm for the unidirectional model. Adding a bidirectional graph convolution structure achieves the lowest MPJPE of 42.9 mm. Additionally, we increased the depth of the bidirectional VIM (denoted as $\dagger$}) to achieve a similar computational cost to the bidirectional graph convolution. The better results (43.2mm vs. 42.9mm) demonstrate that the effectiveness of the GEM module cannot be attributed solely to the increase in computational cost. The inclusion of the GEM module demonstrates its effectiveness in helping the Mamba model learn the local relationships of joints.
@@ V Conclusion
In this paper, we propose a Mamba-Driven Topology Fusion framework for monocular 3D human pose estimation. We first design a bone aware module to estimate the direction and position of bone vectors, providing topological guidance for the Mamba model in capturing dependencies within joint sequences. Based on the Vision Mamba block, we further introduce GEM, which incorporates internal forward and backward Graph Convolutional Networks to enhance Mamba’s ability to capture local joint relationships. Finally, we propose a bone-joint fusion embedding to integrate joint and bone features across different coordinate systems, along with a spatiotemporal refinement module for accurate 3D human pose regression. Our method achieves superior results on the Human3.6M and MPI-INF-3DHP datasets compared to previous methods, achieving a win-win in both speed and accuracy.